# Variability study of classical supergiant X-ray binary 4U 1907+09 using NuSTAR

### v2

@@ Abstract
We investigate the X-ray variability of the supergiant X-ray binary 4U 1907+09 using the new NuSTAR observation of 2024. The source had a relatively stable flux level during previous NuSTAR observations, but the flux varied significantly during the current one. The light curve exhibits dips (off-state) and flares (on-state). The phase-coherent timing analysis during the on-state yields a pulse period of $443.99(4)~{}\mathrm{s}$, showing the pulsar’s continued spin-down. The pulse profiles show an asymmetric double-peaked structure with a phase separation of 0.47 between the two peaks. A cyclotron resonance scattering feature (CRSF) is also detected at $\sim 17.6~{}\mathrm{keV}$, along with its harmonic at $\sim 38~{}\mathrm{keV}$, persisting across all flux states. Flux-resolved spectroscopy reveals that the CRSF remains constant despite a 25-fold change in flux. The spectral parameters like photon index and e-fold energy are out of phase with the pulse shape, whereas cutoff energy is in phase with the pulse shape. The source’s luminosity during the on-state is $2.85\times 10^{35}~{}\mathrm{erg~{}s^{-1}}$, consistent with a “pencil” beam radiation pattern expected at this flux level from a collisionless gas-mediated shock. These results offer further insights into the accretion dynamics and magnetic field geometry of this system.
@@ 1 Introduction
Supergiant X-ray binaries (sgXBs) are binary systems that consist of a compact object and a massive O/B supergiant star. The accretion of stellar wind from the supergiant star onto a compact object is responsible for the X-ray emission in these systems. There are two types of X-ray binaries: classical supergiant high-mass binaries and supergiant fast X-ray transients (SFXTs) (Pradhan et al., 2018). 4U 1907+09 is a classical supergiant high-mass X-ray binary identified in the third Uhuru survey (Giacconi et al., 1971). The pulsar orbiting companion with an eccentricity $e\sim 0.28$ and an orbital period of $P\sim 8.3753$ d (in ’t Zand et al., 1998). Optical and infrared studies suggest an O8-O9 Ia type supergiant donor (Cox et al., 2005; Nespoli et al., 2008). The literature reports a source distance of 5 kpc (Cox et al., 2005; Nespoli et al., 2008). However, the distance to the source calculated by Gaia EDR3 is around 1.9 kpc (Bailer-Jones et al., 2021).
@@
X-ray pulsation with a period of 437.5 s was initially reported by Makishima et al. (1984) utilizing Tenma observations. Early observations of the source revealed a relatively constant spin-down rate (Baykal et al., 2006); however, further observations revealed several torque reversals and spin-down rate variations (Fritz et al., 2006; Inam et al., 2009). INTEGRAL and RXTE measurements show short-term fluctuations in the pulse period compared to long-term changes in spin period rates, consistent with the random walk model (Şahiner et al., 2012). The most recent studies reported spin down of the pulsar (Varun et al., 2019; Tobrej et al., 2023).
@@
4U 1907+09 is a variable X-ray source that exhibits irregular flaring and dipping activities. During the dips, in ’t Zand et al. (1997) did not detect X-ray pulsed emission from 4U 1907+09 using RXTE observations. The typical duration of the dips ranges from a few minutes to 1.5 hours. They suggested that cessation of accretion from the inhomogeneous wind of the companion star may be responsible for dips. Doroshenko et al. (2012) investigated the dipping activity using Suzaku observations and found the pulsations during the off state. The transition between off-state (dips) and on-state (normal flux state) was explained within the framework of “gated accretion”.
@@
The X-ray spectra of 4U 1907+09 are characterized by a power law and an exponential cut-off Schwartz et al. (1980); Marshall and Ricketts (1980); Makishima et al. (1984); Cook and Page (1987); Chitnis et al. (1993); Roberts et al. (2001); Coburn et al. (2002); Baykal et al. (2006); Fritz et al. (2006); Varun et al. (2019); Tobrej et al. (2023). During the binary orbit, inhomogeneous accretion via intense stellar wind caused highly variable hydrogen column density ($n_{H}$) ranging from $1\times 10^{22}$ to $9\times 10^{22}$ cm-2 in ’t Zand et al. (1997). Moreover, the spectra showed a narrow spectral line at around 6.4 keV, which is the Fe $\mathrm{K}\alpha$ emission caused by the fluorescence of materials near the pulsar. Cyclotron-resonant scattering features (CRSFs) at $\sim 19$ keV with a second harmonic at $\sim 39$ keV have been seen using Ginga (Mihara, 1995) and BeppoSAX observations (Cusumano et al., 1998). This implies a surface magnetic field strength of $2.1\times 10^{12}$ G (Cusumano et al., 1998). The flaring activity observed by AstroSat indicates that the CRSF parameters have not been altered (Varun et al., 2019). The existence of cyclotron lines at $\sim 18$ keV and $\sim 38$ keV were recently verified by Tobrej et al. (2023) utilizing NuSTAR. Furthermore, an absorption line possibly due to Ni $\mathrm{K}\alpha$ and Ni $\mathrm{K}\beta$ at $\sim 8$ keV was also noticed.
@@ 2 Observation and data reduction
Figure 1: NuSTAR light curve of 4U 1907+09 in 3-79 keV energy range for FPMA. The grey-shaded region shows the “off-state”. The blue and green-shaded region shows the “on-state” with low and high flux, respectively. (see section  3)
@@
Figure 3: left panel: Pulse profile of 4U 1907+09 in 3-79 keV energy range using FPMA. Right panel: The pulse profile of 4U 1907+09 in high and low flux on-state. The green color corresponds to the low flux state and the blue color corresponds to the high flux state. (sec section 3.1)
@@ 3 Results
During this NuSTAR observation, the source was found to have different X-ray fluxes. We divided the data into off-state and on-state. The effective on-state and off-state exposure are 4.5 ks and 16.5 ks, respectively. Further, we divided the on state into high flux and low flux states. The effective exposure during high flux and low flux in the on state are 4.1 ks and 12.4 ks, respectively. The light curve and different states are shown in Figure 1. First, we performed temporal and spectral analyses on the averaged on-state.
@@ 3.1 Timing analysis
Figure 5: The energy dependence of pulse fraction. The dashed red line corresponds to the pulse fraction in the 3-79 keV energy range. (see section 3.1)
@@ 3.2.1 Averaged on state spectral analysis
We used XSPEC V12.14.0h (Arnaud, 1996) for spectral modeling of the data. A power-law model along with an energy cutoff and smoothing function, is used to represent the continuum spectra observed in X-ray pulsars. We fitted the data by model: const*tbabs(highecut*pow). The model const cross-calibrates instruments FPMA and FPMB. tbabs is Tuebingen-Boulder interstellar medium absorption model. pow is the photon power-law model. highecut is the high energy cutoff model given by,
@@
where $E_{c}$ is cutoff energy (keV) and $E_{f}$ is e-folding energy (keV). The data does not effectively constrain the column density. So, we fixed the $n_{H}$ value at $1.5\times 10^{22}\;\mathrm{cm^{-2}}$ (Maitra and Paul, 2013; Varun et al., 2019). The $\chi^{2}/\mathrm{dof}$ comes out to be $\sim 2202/1119$. An iron line and absorption features around 18 keV and 35 keV due to cyclotron absorption are visible in the residual. So, we added model gauss for the Fe feature and a multiplicative model gabs. So, we added model gauss for the Fe feature and a multiplicative model gabs. The model becomes const*tbabs(gauss+highecut*pow)*gabs*gabs. The $\chi^{2}/\mathrm{dof}$ improved to $\sim 1277/1110$. The best fitting gives the $\Gamma\sim 1.12$ and $E_{\mathrm{cutoff}}\sim 11.8$ keV and $E_{\mathrm{fold}}\sim 12.8$ keV. The Gaussian line is fitted at 6.31 keV with an equivalent width equal to 46 eV, and the absorption lines came out at 17.6 keV and 38.02 keV. The flux came out to be $6.59_{-0.04}^{+0.04}\times 10^{-10}\;\mathrm{erg/cm^{2}/s}$.
@@
To assess the importance of the absorption features in the spectra, we used the simftest script from xspec. It uses Monte Carlo simulations to produce simulated data sets from the original data, which are then used to estimate $\Delta\chi^{2}$ for gabs components. We conducted $10^{4}$ simulations for each gabs component. The $\Delta\chi^{2}$ obtained from simulated data sets is considerably low compared to the observed value for each gabs component. It confirms the significant detection of absorption features at $\sim 17.6$ keV and $\sim 38.02$ keV.
@@
Then we fitted the data by model const*tbabs(gauss+compTT)*gabs*gabs. compTT (Titarchuk, 1994) is an analytic model describing the Comptonization of soft photons in hot plasma and is suitable for describing the continuum shape from X-ray Pulsars. The seed soft photon temperature, plasma temperature, and plasma optical depth are $\sim 0.52$ keV, $\sim 7.87$ keV, and $\sim 6.26$, respectively. The $\chi^{2}/\mathrm{dof}$ for this model is $\sim 1285/1110$ which is slightly higher than the previous model.
@@
Finally, we employed the compmag (Farinelli et al., 2012) model to fit the continuum spectra. The compmag model describes the spectral formation in the accretion column onto the polar cap of a magnetized neutron star, taking into account both thermal and bulk Comptonization processes. Since the magnetic field of 4U 1907+09 is $2\times 10^{12}$ G, the compmag model is suitable for this system. The model becomes const*tbabs(gauss+bbodyrad+compmag)*gabs*gabs. During the fitting, we consider that the accretion velocity increases towards the NS (by fixing betaflag=1) with the index of velocity profile $\eta=0.5$ and terminal velocity at the NS surface $\beta_{0}=0.05$. We fixed the accretion column $r_{0}=0.25$ (in units of the NS Schwartzchild radius) and NS albedo A = 1. Since the norm of compmag is given by $R^{2}_{km}/D^{2}_{10}$, where $R_{km}$ is the radius of the accretion column in km and $D_{10}$ is the distance of the source in units of 10 kpc. Considering the radius of the accretion column is 1 km and the source distance is equal to 1.9 kpc. We fixed the norm at 28. The temperature of seed blackbody temperature ($kT_{bb}$) of compmag is tied with the blackbody temperature of bbodyrad. The remaining parameters of this spectral model were kept free. The $\chi^{2}/\mathrm{dof}$ for this model is $1236/1110$. The optical depth of the accretion column comes out to be $\sim 1.75$. Since the compmag model is applicable for $\tau>1$, we did not further use this model. Also, the low luminosity of the source disfavors the formation of the accretion column. The spectral parameters for different models used in this work are given in Table 1. The NuSTAR spectra with residuals of different models used in this work are shown in Figure 6.
@@ 3.2.3 Flux resolved spectral analysis
As shown in Figure 1, 4U 1907+09 shows the flux variation in the light curve. The spectral data for the off-state, low-flux on-state, and high-flux on-state is shown in Figure 8. We fitted the off-state and high-flux on-state data with the model const*tbabs(bbodyrad+highecut*pow)*gabs. The low flux on-state data required additional models gabs and gauss. We fitted the low flux on-state data with the model. const*tbabs(gauss+bbodyrad+highecut*pow)*
@@
. The $\Gamma$ for off-state, high flux on-state, and low flux on-state come out to be $\sim 1.43$, $\sim 1.12$, and $\sim 1.08$, respectively. At various flux levels, the cyclotron line is found at $\sim 18$ keV. The line width of CRSF during high flux on-state and low flux on-state are $\sim 3.07$ and $\sim 2.5$, respectively. Compared to low flux on state, flux increased 3 times during high flux on state, and flux decreased 8 times during off-state. The e-folding energy during the high flux on-state and low flux on-state are $\sim 4.6$ and $\sim 6.27$, respectively. The spectral parameters for each data are given in Table 2.
@@
$kT_{\mathrm{e}}$ (keV) - $7.87_{-1.38}^{+2.50}$
$\mathrm{\tau}$ $6.26_{-0.17}^{+0.27}$
norm (×$\times 10^{-2}$) - $1.39_{-0.08}^{+0.21}$
compmag $kT_{\mathrm{e}}$ (keV) - - $3.40_{-0.10}^{+0.15}$
$\mathrm{\tau}$ $1.75_{-0.14}^{+0.19}$
@@
gabs1 $E_{\mathrm{CRSF}}$(keV) $17.60_{-0.24}^{+0.12}$ $17.95_{-0.14}^{+0.16}$ $17.60_{-0.24}^{+0.30}$
$\sigma_{\mathrm{CRSF}}$ (keV) $2.23_{-0.36}^{+0.47}$ $2.88_{-0.13}^{+0.15}$ $2.68_{-0.27}^{+0.42}$
strength $1.78_{-0.36}^{+0.47}$ $4.47_{-0.36}^{+0.51}$ $3.06_{-0.52}^{+1.19}$
gabs2 $E_{\mathrm{CRSF}}$(keV) $38.02_{-0.82}^{+0.89}$ $39.55_{-1.29}^{+0.72}$ $34.93_{-0.66}^{+0.94}$
$\sigma_{\mathrm{CRSF}}$ (keV) $5.88_{-0.80}^{+0.98}$ $10.00_{-1.10}^{+1.33}$ $4.04_{-0.82}^{+1.67}$
strength $28.44_{-5.97}^{+8.88}$ $84.41_{-25.15}^{+8.89}$ $11.20_{-2.94}^{+7.12}$
@@
highecut $E_{\mathrm{cutoff}}$(keV) $22.9_{-2.5}^{+1.1}$ $19.3_{-0.78}^{+0.51}$ $18.26_{-0.50}^{+0.41}$
$E_{\mathrm{fold}}$ (keV) $4.6_{-1.5}^{+2.8}$ $4.6_{-0.25}^{+0.37}$ $6.27_{-0.42}^{+0.56}$
gabs1 $E_{\mathrm{CRSF}}$(keV) $18.56_{-0.61}^{+0.64}$ $18.31_{-0.43}^{+0.31}$ $18.09_{-0.23}^{+0.24}$
$\sigma_{\mathrm{CRSF}}$ (keV) $2.20_{-0.44}^{+0.53}$ $3.07_{-0.24}^{+0.19}$ $2.50_{-0.13}^{+0.15}$
strength $3.43_{-0.99}^{+1.10}$ $6.64_{-0.84}^{+0.67}$ $5.05_{-0.43}^{+0.45}$
gabs2 $E_{\mathrm{CRSF}}$(keV) - - $33.69_{-0.83}^{+0.95}$
$\sigma_{\mathrm{CRSF}}$ (keV) - - $2.99_{-0.76}^{+1.28}$
strength - - $5.68_{-2.00}^{+3.62}$

### v3

@@ Abstract
We investigate the X-ray variability of the supergiant X-ray binary 4U 1907+09 using the new NuSTAR observation of 2024. Unlike the previous NuSTAR observations, the source shows significant flux variation during the current one. The light curve exhibits dips (off-state) and flares (on-state). The phase-coherent timing analysis during the on-state yields a pulse period of $443.99(4)~{}\mathrm{s}$, showing the pulsar’s continued spin-down. The pulse profiles show an asymmetric double-peaked structure with a phase separation of 0.47 between the two peaks. A cyclotron resonance scattering feature (CRSF) is also detected at $\sim 17.6~{}\mathrm{keV}$, along with its harmonic at $\sim 38~{}\mathrm{keV}$, persisting across all flux states. Flux-resolved spectroscopy reveals that the CRSF energy remains constant despite a 25-fold change in flux. The spectral parameters, like photon index and e-fold energy, are out of phase with the pulse shape, whereas the cutoff energy is in phase with the pulse shape. The source’s luminosity during the on-state is $2.85\times 10^{35}~{}\mathrm{erg~{}s^{-1}}$, consistent with a “pencil” beam radiation pattern expected at this flux level from a collisionless gas-mediated shock. These results offer further insights into the accretion dynamics and magnetic field geometry of this system.
@@ 1 Introduction
Supergiant X-ray binaries (sgXBs) are binary systems that consist of a compact object and a massive O/B supergiant star. In most cases, the accretion of stellar wind from the supergiant star onto a compact object is responsible for the X-ray emission in these systems. There are two types of X-ray binaries: classical supergiant high-mass X-ray binaries (HMXRBs) and supergiant fast X-ray transients (SFXTs) (Pradhan et al., 2018). The classical supergiant HMXRBs are persistent, with luminosity above $\sim 10^{35}\;\mathrm{erg\;s^{-1}}$. These systems were seen to vary by a factor of $10-100$ (Walter et al., 2015; Kretschmar et al., 2019). SFXTs have low luminosity ($10^{33}-10^{34}\;\mathrm{erg\;s^{-1}}$) and produce short and strong flares with typical lifetimes of hours, reaching luminosity $10^{35}-10^{37}\;\mathrm{erg\;s^{-1}}$ (Sidoli and Paizis, 2018; Kretschmar et al., 2019).
@@
4U 1907+09 is a classical supergiant high-mass X-ray binary identified in the third Uhuru survey (Giacconi et al., 1971). The pulsar orbiting companion with an eccentricity $e\sim 0.28$ and an orbital period of $P\sim 8.3753$ d (in ’t Zand et al., 1998). Optical and infrared studies suggest an O8-O9 Ia type supergiant donor (Cox et al., 2005; Nespoli et al., 2008). The literature reports a source distance of 5 kpc (Cox et al., 2005; Nespoli et al., 2008). However, the distance to the source calculated by Gaia EDR3 is around 1.9 kpc (Bailer-Jones et al., 2021). X-ray pulsation with a period of 437.5 s was initially reported by Makishima et al. (1984) utilizing Tenma observations. Early observations of the source revealed a relatively constant spin-down rate (Baykal et al., 2006); however, further observations revealed several torque reversals and spin-down rate variations (Fritz et al., 2006; Inam et al., 2009). INTEGRAL and RXTE measurements show short-term fluctuations in the pulse period compared to long-term changes in spin period rates, consistent with the random walk model (Şahiner et al., 2012). The most recent studies reported spin down of the pulsar (Varun et al., 2019; Tobrej et al., 2023).
@@
4U 1907+09 is a variable X-ray source that exhibits irregular flaring and dipping activities. During the dips, in ’t Zand et al. (1997) did not detect X-ray pulsed emission from 4U 1907+09 using RXTE observations. The typical duration of the dips ranges from a few minutes to 1.5 hours. They suggested that cessation of accretion from the inhomogeneous wind of the companion star may be responsible for dips. Doroshenko et al. (2012) investigated the dipping activity using Suzaku observations and found the pulsations during the off state. Bozzo et al. (2008) reviewed the theory of wind accretion in HMXRBs and discussed it in the context of SFXT. Moreover, this model was employed by Doroshenko et al. (2012) to explain the transition between off-state (dips) and on-state (normal flux state) in 4U 1907+09. Shakura et al. (2012) proposed a theoretical model for quasi-spherical subsonic accretion onto slowly rotating pulsars. The accreting matter forms a hot quasi-static shell above the magnetosphere. This model was further used to explain the off states in Vela X-1, 4U 1907+09, and GX 301-2 (Shakura et al., 2013).
@@
The X-ray spectra of 4U 1907+09 are characterized by a power law and an exponential cut-off Schwartz et al. (1980); Marshall and Ricketts (1980); Makishima et al. (1984); Cook and Page (1987); Chitnis et al. (1993); Roberts et al. (2001); Coburn et al. (2002); Baykal et al. (2006); Fritz et al. (2006); Varun et al. (2019); Tobrej et al. (2023). During the binary orbit, inhomogeneous accretion via intense stellar wind caused highly variable hydrogen column density ($n_{H}$) ranging from $1\times 10^{22}$ to $9\times 10^{22}$ cm-2 in ’t Zand et al. (1997). Moreover, the spectra showed a narrow spectral line at around 6.4 keV, which is the Fe $\mathrm{K}\alpha$ fluorescence emission caused by the X-ray reflection of primary X-ray emission from the accretion disk or stellar wind. Cyclotron-resonant scattering features (CRSFs) at $\sim 19$ keV with a second harmonic at $\sim 39$ keV have been seen using Ginga (Mihara, 1995) and BeppoSAX observations (Cusumano et al., 1998). This implies a surface magnetic field strength of $2.1\times 10^{12}$ G (Cusumano et al., 1998). The flaring activity observed by AstroSat indicates that the CRSF parameters have not been altered (Varun et al., 2019). The existence of cyclotron lines at $\sim 18$ keV and $\sim 38$ keV was recently verified by Tobrej et al. (2023) utilizing NuSTAR. Furthermore, an absorption line possibly due to Ni $\mathrm{K}\alpha$ and Ni $\mathrm{K}\beta$ at $\sim 8$ keV was also noticed.
@@ 2 Observation and data reduction
Figure 1: NuSTAR light curve (3-79 keV) and hardness ratio (7-30 keV/ 3-7 keV) of 4U 1907+09 for FPMA with a time bin size of 60 sec is shown in upper and lower panels, respectively. The grey-shaded region shows the “off-state”. The blue and green-shaded region shows the “on-state” with low and high flux, respectively. The data corresponding to the unshaded initial portion of lightcurve ( 900 sec) is not used in this work. (see section  3)
@@
Figure 3: Left panel: Pulse profile of 4U 1907+09 in 3-79 keV energy range using FPMA. Right panel: The pulse profile of 4U 1907+09 in high and low flux on-state. The green color corresponds to the low flux state, and the blue color corresponds to the high flux state. (sec section 3.1)
@@ 3 Results
During this NuSTAR observation, the source was found to have different X-ray fluxes. We divided the data into off-state and on-state. The effective on-state and off-state exposures are 4.5 ks and 16.5 ks, respectively. Further, we divided the on state into high-flux and low-flux states. The effective exposure during high flux and low flux in the on state are 4.1 ks and 12.4 ks, respectively. The light curve, Hardness ratio and different states are shown in Figure 1. First, we performed temporal and spectral analyses on the averaged on-state.
@@ 3.1 Timing analysis
Figure 5: The energy dependence of pulsed fraction. The dashed red line corresponds to the pulse fraction in the 3-79 keV energy range. (see section 3.1)
@@ 3.2.1 Averaged on state spectral analysis
We used XSPEC V12.14.0h (Arnaud, 1996) for spectral modeling of the data. A power-law model, along with an energy cutoff and smoothing function, is used to represent the continuum spectra observed in X-ray pulsars. We fitted the data by model: const*tbabs(highecut*pow). The model const cross-calibrates instruments FPMA and FPMB. tbabs is the Tuebingen-Boulder interstellar medium absorption model. pow is the photon power-law model. highecut is the high-energy cutoff model given by
@@
where $E_{c}$ is cutoff energy (keV) and $E_{f}$ is e-folding energy (keV). An artificial discontinuity can be observed at the cutoff energy, and in those cases it is smoothed by including a Gaussian absorption line tied to the cutoff energy. We checked for this feature in our observations and, due to the absence of any aforementioned discontinuity, we did not use the additional Gaussian component during the fitting. The data does not effectively constrain the column density. So, we fixed the $n_{H}$ value at $1.5\times 10^{22}\;\mathrm{cm^{-2}}$ (Maitra and Paul, 2013; Varun et al., 2019). The $\chi^{2}/\mathrm{dof}$ came out to be $\sim 2202/1119$. An iron line and absorption features around 18 keV and 35 keV due to cyclotron absorption are visible in the residual. So, we added model gauss for the Fe feature and a multiplicative model gabs. The model becomes const*tbabs(gauss+highecut*pow)*gabs*gabs. The $\chi^{2}/\mathrm{dof}$ improved to $\sim 1277/1110$. The best fit yielded $\Gamma\sim 1.12$ and $E_{\mathrm{cutoff}}\sim 11.8$ keV, and $E_{\mathrm{fold}}\sim 12.8$ keV. The Gaussian line is fitted at 6.31 keV with an equivalent width equal to 46 eV, and the absorption lines came out at 17.6 keV and 38.02 keV. The flux came out to be $6.59_{-0.04}^{+0.04}\times 10^{-10}\;\mathrm{erg/cm^{2}/s}$.
@@
To assess the importance of the absorption features in the spectra, we used the simftest script from xspec. It uses Monte Carlo simulations to produce simulated data sets from the original data, which are then used to estimate $\Delta\chi^{2}$ for gabs components. We conducted $10^{4}$ simulations for each gabs component. The difference between observed $\chi^{2}$ and maximum $\chi^{2}$ obtained from simulated data sets for $\sim 17.6$ keV and $\sim 38.02$ keV gabs component are $\sim 400$ and $\sim 54$. It confirms the $>5\sigma$ detection of absorption features at $\sim 17.6$ keV and $\sim 38.02$ keV.
@@
Then we fitted the data by model const*tbabs(gauss+compTT)*gabs*gabs. compTT (Titarchuk, 1994) is an analytic model describing the Comptonization of soft photons in hot plasma and is suitable for describing the continuum shape from X-ray pulsars. The seed soft photon temperature, plasma temperature, and plasma optical depth are $\sim 0.52$ keV, $\sim 7.87$ keV, and $\sim 6.26$, respectively. The $\chi^{2}/\mathrm{dof}$ for this model is $\sim 1285/1110$, which is slightly higher than the previous model.
@@
Finally, we employed the compmag (Farinelli et al., 2012) model to fit the continuum spectra. The compmag model describes the spectral formation in the accretion column onto the polar cap of a magnetized neutron star, taking into account both thermal and bulk Comptonization processes. Since the magnetic field of 4U 1907+09 is $2\times 10^{12}$ G, the compmag model is suitable for this system. The model becomes const*tbabs(gauss+bbodyrad+compmag)*gabs*gabs. During the fitting, we consider that the accretion velocity increases towards the NS (by fixing betaflag=1) with the index of velocity profile $\eta=0.5$ and terminal velocity at the NS surface $\beta_{0}=0.05$. We fixed the accretion column $r_{0}=0.25$ (in units of the NS Schwarzschild radius) and NS albedo A = 1. Since the norm of compmag is given by $R^{2}_{km}/D^{2}_{10}$, where $R_{km}$ is the radius of the accretion column in km and $D_{10}$ is the distance of the source in units of 10 kpc. Considering the radius of the accretion column is 1 km, and the source distance is equal to 1.9 kpc. We fixed the norm at 28. The temperature of the seed blackbody temperature ($kT_{bb}$) of compmag is tied with the blackbody temperature of bbodyrad. The remaining parameters of this spectral model were kept free. The $\chi^{2}/\mathrm{dof}$ for this model is $1236/1110$. The optical depth of the accretion column came out to be $\sim 1.75$. Since the compmag model is applicable for $\tau<1$, we did not further use this model. Also, the low luminosity of the source disfavors the formation of the accretion column. The spectral parameters for different models used in this work are given in Table 1. The NuSTAR spectra with residuals of different models used in this work are shown in Figure 6.
@@ 3.2.3 Flux resolved spectral analysis
As shown in Figure 1, 4U 1907+09 shows the flux variation in the light curve. The spectral data for the off-state, low-flux on-state, and high-flux on-state is shown in Figure 8. We fitted the off-state and high-flux on-state data with the model const*tbabs(highecut*pow)*gabs. The low flux on-state data required additional models gabs and gauss. We fitted the low flux on-state data with the model. const*tbabs(gauss+highecut*pow)*
@@
. The $\Gamma$ for off-state, high flux on-state, and low flux on-state come out to be $\sim 1.43$, $\sim 1.12$, and $\sim 1.08$, respectively. At various flux levels, the cyclotron line is found at $\sim 18$ keV. The line width of CRSF during high flux on-state and low flux on-state is $\sim 3.07$ and $\sim 2.5$, respectively. Compared to low flux on-state, flux increased 3 times during high flux on-state, and flux decreased 8 times during off-state. The e-folding energy during the high flux on-state and low flux on-state is $\sim 4.6$ and $\sim 6.27$, respectively. The spectral parameters for each data are given in Table 2.
@@
$kT_{\mathrm{e}}$ (keV) - $7.9_{-1.4}^{+2.5}$
$\mathrm{\tau}$ $6.3_{-0.2}^{+0.3}$
norm (×$\times 10^{-2}$) - $1.4_{-0.1}^{+0.2}$
compmag $kT_{\mathrm{e}}$ (keV) - - $3.4_{-0.1}^{+0.2}$
$\mathrm{\tau}$ $1.8_{-0.1}^{+0.2}$
@@
gabs1 $E_{\mathrm{CRSF}}$(keV) $17.6_{-0.2}^{+0.1}$ $17.9_{-0.1}^{+0.2}$ $17.6_{-0.2}^{+0.3}$
$\sigma_{\mathrm{CRSF}}$ (keV) $2.2_{-0.4}^{+0.5}$ $2.9_{-0.1}^{+0.2}$ $2.7_{-0.3}^{+0.4}$
strength $1.8_{-0.4}^{+0.5}$ $4.5_{-0.4}^{+0.5}$ $3.1_{-0.5}^{+1.2}$
gabs2 $E_{\mathrm{CRSF}}$(keV) $38.0_{-0.8}^{+0.9}$ $39.6_{-1.3}^{+0.7}$ $34.9_{-0.7}^{+0.9}$
$\sigma_{\mathrm{CRSF}}$ (keV) $5.9_{-0.8}^{+1.0}$ $10.0_{-1.1}^{+1.3}$ $4.0_{-0.8}^{+1.7}$
strength $28.4_{-6.0}^{+8.9}$ $84.4_{-25.2}^{+8.9}$ $11.2_{-2.9}^{+7.1}$
@@
highecut $E_{\mathrm{cutoff}}$(keV) $22.9_{-2.5}^{+1.1}$ $19.3_{-0.8}^{+0.5}$ $18.3_{-0.5}^{+0.4}$
$E_{\mathrm{fold}}$ (keV) $4.6_{-1.5}^{+2.8}$ $4.6_{-0.3}^{+0.4}$ $6.3_{-0.4}^{+0.6}$
gabs1 $E_{\mathrm{CRSF}}$(keV) $18.6_{-0.6}^{+0.6}$ $18.3_{-0.4}^{+0.3}$ $18.1_{-0.2}^{+0.2}$
$\sigma_{\mathrm{CRSF}}$ (keV) $2.2_{-0.4}^{+0.5}$ $3.1_{-0.2}^{+0.2}$ $2.5_{-0.1}^{+0.2}$
strength $3.4_{-1.0}^{+1.1}$ $6.6_{-0.8}^{+0.7}$ $5.1_{-0.4}^{+0.5}$
gabs2 $E_{\mathrm{CRSF}}$(keV) - - $33.7_{-0.8}^{+1.0}$
$\sigma_{\mathrm{CRSF}}$ (keV) - - $3.0_{-0.8}^{+1.3}$
strength - - $5.7_{-2.0}^{+3.6}$